OpenStudy (anonymous):

solve for F c=5/9(F-32)

OpenStudy (anonymous):

\[c= \frac{5}{9f-288}\]

OpenStudy (anonymous):

\[(9f-288)=\frac{5}{c}\]See it now?

OpenStudy (anonymous):

nope i dont see how you got the first step

OpenStudy (anonymous):

Multiplies everything in the brackets by 9 9(f-32)=9f-288

OpenStudy (anonymous):

=9f-9*32

OpenStudy (anonymous):

this is the formula for converting from Fahrenheit to Celsius it is \[C=\frac{5}{9}(F-32)\]

OpenStudy (shane_b):

I think you're over-complicating this. 1) Multiply both sides by 9/5 A 2) Add 32 to each side 3) Simplify

OpenStudy (anonymous):

to go the other way, solve for \(F\)

OpenStudy (anonymous):

what @Shane_B said

OpenStudy (shane_b):

(disregard the A above...typo :)

OpenStudy (shane_b):

Following what I said: \[C=\frac{5}{9}(F-32)\]\[\frac{9}{5}C=F-32\]\[\frac{9}{5}C+32=F\]

OpenStudy (anonymous):

aah,,,ok,, thank you very much

OpenStudy (shane_b):

yw